OpenStudy (anonymous):

the sum of the ages of a son and father is 56 years. after 4 years the age of father will be 3 times of sons.wat is the age of son ?

OpenStudy (anonymous):

44 and 12

OpenStudy (anonymous):

Let, fathers age is x and that of son is 56-x Now, x + 4 = 3 (60 - x) or, x + 3x = 180 - 4 or, 4x = 176 or, x = 176/4 = 44 So, age of son = 56 - 44 = 12

OpenStudy (anonymous):

ok thnxx !! is der any other method ! i got ths method but, i was dng in other sense ?

OpenStudy (anonymous):

let, fathers age is x and son's age is y. Now, x + y = 56 and, x + 4 = 3(y + 4) Now you can solve this two equation to find the ages.. Well this is actually the same as the previous one.

OpenStudy (anonymous):

in options its given as 10,11,12,13... i took all the options at once and verified it from quest.. but i dnt get it, if i take age of son as 12,den 12*3=36 (3 times as given in quest), den 36-4=32 & 12-4=8 -- both the ages before 4 years, so am nt gettin the rite ans i.e. sum of ages before 4 years= 56..

OpenStudy (anonymous):

toriqur ??????

OpenStudy (anonymous):

not man. if you take 12 years for son, after 4 years, he will be 12+4 = 16. Now father is 44. After 4 years he will be 44+4 = 48 that equals to 16*3. Got it?

OpenStudy (anonymous):

ya i fot !!

OpenStudy (anonymous):

got*

OpenStudy (anonymous):

ok.